OpenStudy (anonymous):

What is the function rule for the graph shown? A graph of the ordered pairs: nine comma zero, eleven comma two, twelve comma three, and fourteen comma five. y = x + 9 y = x – 9 y = 9x y equals one over nine x

OpenStudy (anonymous):

Let's use 2 points to get our gradient: (9,0) and (11,2) m = (y2-y1)/(x2-x1) m = (2-0)/(11-9) m = 2/2 m = 1 So we know it is one of the first two. Do you think you can take it from here?